OpenStudy (anonymous):

Factor completely: −5x3 + 10x2 Prime −x(5x2 − 10x) 5x(−x2 + 2x) −5x2(x − 2)

OpenStudy (anonymous):

third one

OpenStudy (primeralph):

prime? me?

OpenStudy (ivettef365):

-5x^3 +10x^2 (-5) and (10) = -5 x^3 and x^2 = x^2 (choose the smallest) therefore you have -5x^2(x-2)